OpenStudy (fifib):

HOW AM I SUPPOSED TO CALCULATE THIS?? Formic acid, HCOOH, which causes the sting felt from some insect bites, is 1.48% dissociated in 0.50 M HCOOH (aq) at a certain temperature. a. Calculate [H+]. b. Calculate Ka for HCOOH at this temperature. c. Determine pH of 0.50 M HCOOH.

OpenStudy (photon336):

\[HCOOH \leftarrow \rightarrow H^+~~HCOO^{-}\] \[K_{a} = \frac{ [HCOO^-][H^+] }{ [HCOOH] }\]

OpenStudy (fifib):

do i have to substitute their molar mass ? @Photon336

OpenStudy (aaronq):

which part are you working on right now?

OpenStudy (fifib):

I have not started @aaronq photon said he was going to come back in a while . I have no clue how to do this

OpenStudy (aaronq):

a) is pretty simple, you know that the acid dissociates like: \(\sf HA\rightarrow H^++A^-\), so just find 1.48% of the initial concentration of the acid (0.5 M) and thats equal to \([H^+]\). for b) you would use the info from a). the concentrations are: \(\sf [H^+]=[A^-]\) and \(\sf [HA]_{eq}=[HA]_0-[H^+]\) for c) use \(\sf pH=-log[H^+]\)

OpenStudy (aaronq):

If you need to visualize b) further, you can make an I.C.E table (initial, change and equilibrium concentrations).

OpenStudy (fifib):

I'm not that bright of a student. i need a simpler explanation. i appreciate your effort :) i just dont know how to do this at all @aaronq

OpenStudy (aaronq):

oh dont say that, you probably just havent really applied yourself :P so for a) it's a simple concept. if i asked you what is 5% of $1.00, what would you say?

OpenStudy (fifib):

.05?

OpenStudy (aaronq):

right! how did you figure that out?

OpenStudy (fifib):

i changed 5% to 5/100 and multiplied it into 1

OpenStudy (aaronq):

exactly. so what is 1.48% of 0.5 M?

OpenStudy (fifib):

.0074

OpenStudy (aaronq):

yup! so \(\sf [H^+]=0.0074~M\)

OpenStudy (fifib):

yay :) what is the formula for b?

OpenStudy (aaronq):

this goes back to dissociation equation: \(\sf HCOOH\rightarrow H^++HCOO^-\)

OpenStudy (aaronq):

what photon wrote it the equilibrium expression for the dissociation. something you would need to know how to write if you were going to the question alone. We need to make an I.C.E. table, we start off with 0.5 M of the acid and none of the products. \( \sf [HCOOH] ~~~~~~ [H^+]~~~~[HCOO^-] \) I 0.5 M 0 0 C E then the change is, 0.0074, which we will represent as x for now. \( \sf [HCOOH] ~~~~~~ [H^+]~~~~[HCOO^-] \) I 0.5 M 0 0 C -x +x +x E finally we equilibrium concentrations are just putting initial and change values together. \( \sf [HCOOH] ~~~~~~ [H^+]~~~~[HCOO^-] \) I 0.5 M 0 0 C -x +x +x E 0.5 -x x x We then use these in the equilibrium expression, try plugging them in.

OpenStudy (fifib):

Whoa

OpenStudy (aaronq):

it's not complicated, i just wrote one step at a time to help you see what you do at each step.

OpenStudy (fifib):

which one do i use though

OpenStudy (aaronq):

the equilibrium (E) values

OpenStudy (fifib):

Um ok

OpenStudy (aaronq):

I'll give you a hand. We have the equilibrium expression, so we plug in the values from E. \(\sf K_{a} = \dfrac{ [HCOO^-][H^+] }{ [HCOOH] }=\dfrac{x*x}{0.5-x}=\dfrac{x^2}{0.5-x}\) but remember we found out what x was in part a)? \(\sf [H^+]=x=0.0074\) \(\sf K_a=\dfrac{x^2}{0.5-x}=\dfrac{(0.0074)^2}{0.5-0.0074}\)

OpenStudy (fifib):

I got such a huge number @aaronq

OpenStudy (aaronq):

it should be a very small number, remember this is the dissociation constant of a weak acid that only dissociates 1.48%.

OpenStudy (fifib):

but i keep getting 0.00011116 what i meant by huge is it has many number not a Big number

OpenStudy (aaronq):

ohh that's actually the right answer! um huge and big are synonyms, btw :P haha try writing that number in scientific notation though, it's usually how it's expressed and you should be able to write it as so.

OpenStudy (fifib):

Lol i feel so dumb atm

OpenStudy (fifib):

so is it 1.1116 × 10-4 @aaronq

OpenStudy (aaronq):

\(\huge \checkmark\)

OpenStudy (fifib):

yes! so that was b right

OpenStudy (aaronq):

yes, that was b)

OpenStudy (aaronq):

for c) use \(\sf pH=-log[H^+]\) you already know \(\sf [H^+]\) from a)

OpenStudy (fifib):

ohhh okay thank you :) :)

OpenStudy (aaronq):

no problm (: